OpenStudy (anonymous):

Find the common ratio for the following sequence. 1/2, -1/4, 1/8, -1/16, ...

OpenStudy (anonymous):

-1/2

OpenStudy (tad1):

-1/2

OpenStudy (anonymous):

You are looking for the fraction you multiply by each term to get the next one.\[ -1/2 \times -1/2 = -1/4\], etc.

OpenStudy (amistre64):

divide the a term by the term before it; and you get the ratio

OpenStudy (amistre64):

1/8 ---- = -4/8 = -1/2 -1/4

OpenStudy (anonymous):

polpak:\(-1/2\times -1/2\neq-1/4\) :P

OpenStudy (anonymous):

is: ((-1)^i+1)(1/2)^i where i= 1, 2, 3 .....

OpenStudy (anonymous):

Bah.. had an extra negative in there.. You get the point ;p

OpenStudy (anonymous):

polpak -1/2 x -1/2= 1/4 - x -= + - x += -

OpenStudy (anonymous):

i=1 => ((-1)^2)(1/2) = 1/2 i=2 => ((-1)^3)(1/4) = -1/4 just reemplace i for any numbre